OpenStudy (anonymous):

find the parametric equation for (x+9)^2 + (y-4)^2 = 49

OpenStudy (anonymous):

In general for parametric equations \[x=x _{0}+acos(t) \] and \[y = y _{0}+asin(t)\] so for your equation x=-9+7cos(t) and y=4+7sin(t). Hope that helps